OpenStudy (anonymous):

Given the velocity v=ds/dt and initial position of a body moving along a coordinate line, find the body's position at time t. v = sin(pi(t)) s(-6)=0.

OpenStudy (anonymous):

Given the velocity v=ds/dt and initial position of a body moving along a coordinate line, find the body's position at time t. v = sin(pi(t)) s(-6)=0.

OpenStudy (anonymous):

\[\int\limits_{a}^{b} v(t) = s(t) = -(1/\pi)*\cos(\pi(t))+C\] s(t)=0 at t=-6 so C=(1/pi) So finally, \[s(t)=[-\cos(\pi*t)+1]/\pi\]

OpenStudy (anonymous):

Thanks.